Suppose A is an n×n matrix, and let v1,.....vn belong to R^n. Suppose {Av1,.....Avn} is linearly independent prove that A is non singular

1
Expert's answer
2021-12-30T10:46:05-0500

"\begin{aligned}
&\text { Given that }\left\{A v_{1}, A v_{2}, A v_{3} \ldots A v_{n}\right\} \text { is linearly } \\
&\text { independent vectors of } \mathbb{R}^{n} \\
&\text { As } \mathbb{R}^{n} \text { has dimension } n \text {, this means that } \\
&\left\{A v_{1}, A v_{2}, A v_{3} \ldots A v_{n}\right\} \text { must be a basis for } \mathbb{R}^{n} \\
&\text { Therefore, the column space of } A \text {, which consists of } \\
&\text { vectors of the form }\left\{A v: v \in \mathbb{R}^{n}\right\} \text { is of dimension } n \text {. } \\
&\text { In other words, rank }(A)=n \\
&\text { We want to show that } A \text { is non-singular } \\
&\text { That is, } x \in \mathbb{R}^{n} \text { and if } A x=0 \Rightarrow x=0 \\
&\text { Let } A x=0 \text { for some } x \in \mathbb{R}^{n} \\
&\text { Thus, we must have } x \in \mathrm{ker}(A) \text { (the null space of the } \\
&\text { given matrix) which is defined as } \\
&\text { ker }(A)=\left\{x \in \mathbb{R}^{n}: A x=0\right\} \text { and } \\
&\text { null }(A)=\operatorname{dim}(\text { ker }(A)) \\
&\text { As } A: \mathbb{R}^{n} \rightarrow \mathbb{R}^{n} \text { by Rank-Nullity theorem, we must } \\
&\text { have } \\
&\text { rank } A+\text { null } A=n \\
&\text { But rank } A=n \text { so that } \\
&n+\text { null } A=n \Rightarrow \text { null } A=0 \\
&\text { Thus, we must have } \\
&\text { ker }(A)=\left\{x \in \mathbb{R}^{n}: A x=0\right\}=\{0\} \text { as this is the } \\
&\text { only subspace having rank } 0 \\
&\text { Therefore, } \\
&x \in \text { ker }(A)=\left\{x \in \mathbb{R}^{n}: A x=0\right\}=\{0\} \Rightarrow x=0 \\
&\text { We have shown that for all } x \in \mathbb{R}^{n}, \text { we have } \\
&A x=0 \Rightarrow x=0 \\
&\text { Therefore, the matrix } A \text { is non-singular. }
\end{aligned}"